# TreeDGS: Aerial Gaussian Splatting for Distant DBH Measurement

### v1

@@ 4.1 Experimental settings
Figure 7: LiDAR intensity vs. TreeDGS opacity as a reliability signal at a same 5m slice height (Plot 2). (a) UAV LiDAR point cloud colored by normalized return intensity. (b) TreeDGS surface samples colored by learned opacity (0–1). Bottom: zoomed top-down views of two example stems ((a-1),(a-2) intensity; (b-1),(b-2) opacity). Compared to LiDAR intensity, the opacity values form a clearer high-confidence core around the stem cross-sections, which we use as a per-point confidence cue for DBH circle fitting. LiDAR and TreeDGS slice have 281.4K and 66.6K points, respectively.
Figure 8: Qualitative comparison of point density and trunk completeness (Plot 1). (a) SfM points are sparse with limited stem coverage. (b) OpenMVS densification improves density but remains incomplete on trunks under canopy occlusion. (c) TreeDGS extracts a dense point set from the optimized Gaussians via opacity-guided sampling and RaDe-GS depth-aware cumulative-opacity tagging, yielding more continuous trunk support near breast height. (Max 100 candidate draws per splat; retained points vary with opacity.) (d) UAV LiDAR provides strong structure but can be still stem-sparse/contaminated depending on scan geometry and occlusion. Colors indicate height above ground.
@@
We report four standard error measures between estimated and field DBH: root mean squared error (RMSE), relative RMSE (RRMSE, normalized by the mean field DBH), mean absolute error (MAE), and mean error (ME, bias). All errors are reported in centimeters, and lower is better for RMSE/RRMSE/MAE while ME closer to 0 indicates lower systematic bias.
@@
Figure 10: Mean-only vs. proposed point sampling from Gaussian Splats. (a) Exporting only Gaussian means [27] (one point per splat) yields an extremely sparse, fragmented cloud that is insufficient for downstream tasks such as 3D segmentation (ForestFormer3D [48]). (b) Our opacity-guided surface sampling densifies geometry and preserves stem surfaces and canopy structure. Colors indicate height.
@@ 4.2 Results
Ablations also highlight that the fitting model must match the reconstructed geometry. Applying a classical cylinder fit [32] to points sampled from TreeDGS performs poorly (13.29 cm RMSE; ME =6.46=-6.46 cm), consistent with our discussion in Section 3.6: TreeDGS sampling produces slice-wise cross-sections rather than a thin ring of surface points. Similarly, unweighted circle fitting is substantially less accurate (9.61 cm RMSE), showing that down-weighting low-confidence samples using multi-view opacity is essential.
@@ 5 Conclusion
We introduced TreeDGS, a UAV RGB-only pipeline that repurposes 3D Gaussian Splatting as an opacity-weighted density sampling method for trunk measurement, rather than only a view-synthesis renderer. Building on SfM/MVS initialization, TreeDGS optimizes a Gaussian field and then converts it into a measurement-ready stem point set via depth-aware cumulative-opacity sampling and multi-view reliability weighting, enabling robust circle fitting at breast height. In doing so, our approach bridges a key gap left by prior UAV LiDAR and UAV-SfM DBH pipelines that operate on inherently sparse or view-limited point clouds [35, 34], and demonstrates that Gaussian splats can support accurate trunk-level metrics from commodity aerial imagery [22].
TreeDGS has important limitations. First, it relies on trunk visibility: the breast-height band must be observed in at least a few frames (in practice, $\sim$2+ informative views for SfM/MVS and Gaussian optimization to constrain the stem surface. When this band is fully occluded in most views, the method cannot recover the missing geometry, regardless of how densely the Gaussian field is sampled; sampling density cannot compensate for geometry that is never sufficiently observed in the input imagery. Second, DBH estimation is sensitive to 3D semantic segmentation quality. Because diameter is fit from trunk-isolated points, false-positive trunk labels from nearby vegetation can bias the estimate. Although multi-view opacity weighting mitigates this effect, the pipeline remains subject to the failure modes of the trunk classifier [48]. This sensitivity can also contribute to poorer LiDAR baseline performance in challenging plots: when breast-height stem returns a small number of vegetation false positives can disproportionately bias cylinder fits.
A promising direction is to reduce dependence on post-processing (sampling and segmentation) by incorporating explicit trunk primitives into Gaussian optimization. Jointly optimizing a small set of parameterized stem elements (e.g., tapered cylinders) together with the Gaussian field would enable DBH to be inferred directly from model parameters and would provide a stronger inductive bias under partial trunk visibility. Such a hybrid representation may also support end-to-end supervision when ground-truth field DBH is available.

### v2

@@ 4.1 Experimental settings
Figure 7: LiDAR intensity vs. TreeDGS opacity as a reliability signal at a same 5m slice height (Plot 2). (a) UAV LiDAR point cloud colored by normalized return intensity. (b) TreeDGS surface samples colored by learned opacity (0–1). Bottom: zoomed top-down views of two example stems ((a-1),(a-2) intensity; (b-1),(b-2) opacity). Compared to LiDAR intensity, the opacity values form a clearer high-confidence core around the stem cross-sections. This provides per-point confidence cue for DBH circle fitting. LiDAR has 281.4K points and TreeDGS has 66.6K points.
Figure 8: Qualitative comparison of point density and trunk completeness (Plot 1). (a) SfM points are sparse with limited stem coverage. (b) OpenMVS densification improves density but remains incomplete on trunks under canopy occlusion. (c) TreeDGS extracts a dense point set from the optimized Gaussians via opacity-guided sampling and RaDe-GS depth-aware cumulative-opacity tagging, yielding more continuous trunk support near breast height. (Max 100 candidate draws per splat; retained points vary with opacity.) (d) UAV LiDAR provides strong structure but can still be stem-sparse/contaminated depending on scan geometry and occlusion. Colors show height above ground.
@@
We report four error measures between estimated and field DBH: root mean squared error (RMSE), relative RMSE (RRMSE, normalized by the mean field DBH), mean absolute error (MAE), and mean error (ME, bias). All errors are in cm, lower is better for RMSE/RRMSE/MAE, and ME closer to 0 indicates lower systematic bias.
@@
Figure 10: Mean-only vs. proposed point sampling from Gaussian Splats. (a) Exporting only Gaussian means [27] (one point per splat) yields an extremely sparse, fragmented cloud that is insufficient for downstream tasks such as 3D segmentation (ForestFormer3D [48]). (b) Our opacity-guided surface sampling densifies geometry and preserves stem surfaces and canopy structure.
@@ 4.2 Results
Ablations also show that the fitting model must match the reconstructed geometry. Applying a classical cylinder fit [32] to points sampled from TreeDGS performs poorly (13.29 cm RMSE; ME =6.46=-6.46 cm), consistent with our discussion in Section 3.6: TreeDGS sampling produces slice-wise cross-sections rather than a thin ring of surface points. Similarly, unweighted circle fitting is substantially less accurate (9.61 cm RMSE), showing that down-weighting low-confidence samples using multi-view opacity is a better approach.
@@ 5 Conclusion
We introduced TreeDGS: a UAV RGB-only pipeline that repurposes 3D Gaussian Splatting as an opacity-weighted density sampling method for trunk measurement, rather than only a view-synthesis renderer. Building on SfM/MVS initialization, TreeDGS optimizes a Gaussian field and then converts it into a measurement-ready stem point set via depth-aware cumulative-opacity sampling and multi-view reliability weighting, enabling robust circle fitting at breast height. This approach bridges a key gap left by prior UAV LiDAR and UAV-SfM DBH pipelines that operate on inherently sparse or view-limited point clouds [35, 34], and demonstrates that Gaussian splats can support accurate trunk-level metrics from commodity aerial imagery [22].
TreeDGS has important limitations. First, it relies on trunk visibility: the breast-height band must be observed in $\sim$2–3+ informative views for SfM/MVS and Gaussian optimization to constrain the stem surface. When this band is fully occluded in most views, the method cannot recover the missing geometry, regardless of how densely the Gaussian field is sampled; sampling density cannot compensate for geometry that is never sufficiently observed in the input imagery. Second, DBH estimation is sensitive to 3D semantic segmentation quality. Because diameter is fit from trunk-isolated points, false-positive trunk labels from nearby vegetation can bias the estimate. Although multi-view opacity weighting mitigates this effect, the pipeline remains subject to the failure modes of the trunk classifier [48]. This sensitivity can also contribute to poorer LiDAR baseline performance in challenging plots: breast-height stem returns with a small number of vegetation false positives can disproportionately bias cylinder fits.
A promising direction is to reduce dependence on post-processing (sampling and segmentation) by incorporating explicit trunk primitives into Gaussian optimization. Jointly optimizing a small set of parameterized stem elements (e.g., tapered cylinders) together with the Gaussian field would enable DBH to be inferred directly from model parameters and would provide a stronger inductive bias under partial trunk visibility. Such a hybrid representation may also support end-to-end supervision when ground-truth field DBH is available for future learning-based systems.